# Finite-time convergence to an $\epsilon$-efficient Nash equilibrium in potential games

### v2

@@ 4.1 Reduced feedback
Log-linear learning requires players to access their utilities for all possible actions given the other players’ actions. Having such full-information feedback when action sets are large can be demanding. Binary log-linear learning [1, 20] alleviates this limitation by requiring two-point feedback, reducing the feedback needed by a factor $A$ per round. Now, we briefly review the binary-log-linear learning rule. Then, we derive the first finite-time convergence bound of binary log-linear learning to an $\epsilon$-efficient Nash equilibrium, showing that the deterioration in convergence time is a constant.
@@
$t\geq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)\approx\tilde{\Omega}\left(N^{2}A^{5}\left(\frac{A^{N}}{\epsilon}\right)^{\frac{N}{\max\{\epsilon,\Delta\}}}\right).$
Theorem 2 shows the first finite-time convergence guarantee for binary log-linear learning, we provide its detailed proof in Appendix C. It is remarkable that with significantly less feedback per round, binary log-linear achieves the same convergence speed as log-linear learning up to a factor of $8$ (Theorem 1). This raises the question of whether two-point feedback is sufficient for learning $\epsilon$-efficient Nash equilibria or whether the convergence time bounds we proved for log-linear learning are loose.
@@ 4.2 Perturbed log-linear learning
Classical log-linear learning relies on two limiting assumptions: 1) Players have access to their exact utilities. However, in real-world applications, the presence of noise is typical as uncertainties and hidden factors generate inexact measurements. 2) Players are rational. However, empirical evidence suggests that players have limited rationality and therefore may occasionally deviate from the log-linear learning rule in practical scenarios. Our next result generalizes Theorem 1 to the case where the log-linear learning rule is subject to small perturbations. As we will show, this generalization can address the two limitations above.
@@
We provide a proof in Appendix D.4. Compared to the result of [34], we considerably improve the Lipschitz constant by using the mixing time bound based on log-Sobolev inequalities (Lemmas 1 and 3). In particular, [34, Lemma 24] entails a Lipschitz constant $L=\tilde{\mathcal{O}}((e/p_{\min})^{N})$ while Lemma 4 implies that $L=\tilde{\mathcal{O}}(1/(\mu_{\min}p_{\min}^{3}))$. [^4] This can be seen by injecting the log-Sobolev bound of lemma 3 into our Lemma above. Now, we state the main result of this section.
Theorem 3.
@@
Theorem 3 shows that small perturbations of the log-linear learning rule do not compromise the convergence to an $\epsilon$-efficient NE. In particular, if the players follow a learning rule $P$ with $\|P-P_{\ell}\|_{2}=\mathcal{O}(\epsilon/(L\sqrt{A^{N}}))$ then they converge to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$. On the other hand, due to the unavailability of the stationary distribution of the perturbed learning rule, we suffer an extra factor of $(N/p_{\min})^{N}/N!$ in the convergence time guarantee compared to log-linear learning.
@@ 4.2.1 Corrupted utilities with additive noise
where $\xi_{i}(a_{i},a_{-i})\in[-\xi,\xi]$ is a bounded noise term. Alternatively, the noise could be assumed to be centered i.i.d. random variables with bounded variance [14].
Using Theorem 3, we hereafter prove that log-linear learning is robust to noisy feedback.
Corollary 2.
Consider the setting of Theorem 3 with noise-corrupted utilities as in Equation (9). If all players adhere to log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$ and $\xi\leq 1/(2\beta)$, then
@@
for $t=\Omega\left(N^{3/2}A^{3}e^{N+\beta(1+2\xi)(N+3)}\log\frac{1}{\epsilon^{2}}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
The proof is provided in Appendix D.2. Corollary 2 shows that log-linear learning with corrupted utilities converges to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$ if the corruption magnitude $\xi$ is sufficiently small. Our finite-time convergence result extends previous works on robust learning which provide asymptotic guarantees [14, 16, 6]. The key to this result lies in showing that the transition matrix of the Markov chain induced by corrupted utilities is close to its corruption-free counterpart.
@@ 4.2.2 Log-linear learning mixed with uniform exploration
In the following, we assume players occasionally explore actions randomly. A modification of log-linear learning based on the fixed-share algorithm [12] can reflect such a random behavior. In the so-called fixed-share log-linear learning, a player $i$ is randomly chosen and allowed to alter her action. Player $i$ samples her new action from the following distribution:
@@
The exploration parameter $\xi\in(0,1)$ determines how likely a player is to act randomly, where a value of $\xi=1$ corresponds to a uniform action sampling while $\xi=0$ corresponds to standard log-linear learning. For simplicity, we focus on the full-information case but fixed-share log-linear learning can easily be adapted to the binary setting. Note that this modification resembles the $\epsilon$-Hedge strategy [11] in the expert advice literature, and under binary feedback, this modification resembles the Epx3.P strategy [3, 7] in the multi-armed bandit literature. Here, the fixed share $\xi/A$ ensures a lower bound on the exploration.
Without explicit knowledge of the stationary distribution of this learning rule, we can apply Theorem 3 to deduce the following result.
Corollary 3.
Consider the setting of Theorem 3, where all players adhere to fixed-share log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then, for any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$ we have:
@@
for $t=\Omega\left(N^{3/2}A^{N+3}e^{\beta(N+3)}/(1-\xi)^{N+3}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
We provide a proof in Appendix D.3. Corollary 3 guarantees the convergence of fixed-share log-linear learning to an $\epsilon$-efficient Nash equilibrium in time polynomial in $1/\epsilon$ if the exploration parameter $\xi$ is sufficiently small. The key to this result is to show that the transition matrix of fixed-share log-linear learning is close to the transition matrix of the unperturbed learning rule in terms of the $\ell_{2}$ distance.
@@ Appendix A Markov chains and mixing times
In the following, we define the log-Sobolev constant and the spectral gap of a Markov chain. We then discuss the implications of Lemma 1 and provide a proof of the lemma.
Definition 2 ([10]).
@@
As mentioned in Section 2.2, using the log-Sobolev constant $\rho(PP^{*})$ can often significantly improve classical mixing time bounds based on the spectral gap $\lambda(PP^{*})$. Such classical bounds are of the form [26]:
@@
where $C$ is some constant. Diaconis and Saloff-Coste [10] in Lemma $3.1$ show that the log-Sobolev constant $\rho(PP^{*})$ is upper-bounded by the spectral gap $\lambda(PP^{*})$ as follows: $2\rho(PP^{*})\leq\lambda(PP^{*})$. Thus, roughly speaking if
@@
Proof of Lemma 1.
Let the relative entropy be defined as $D(\mu^{t}:\mu):=\sum_{a\in\mathcal{A}^{N}}\mu^{t}(a)\log\frac{\mu^{t}(a)}{\mu(a)}$. Then, for a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ with irreducible transition matrix $P$ the relative entropy $D(\mu^{t}:\mu)$ decays at the following rate [21]:
@@
Note that $\rho(PP^{*})<1$ since $2\rho(PP^{*})\leq\lambda(PP^{*})$ by Lemma 3.1 in [10] and for the spectral gap $\lambda(PP^{*})$ it is known that $\lambda(PP^{*})<1$ [15]. To ensure that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon$, we derive the following lower bound on $t$:
@@ Appendix B Convergence of log-linear learning
We first prove Lemma 2 and then state a novel lemma which we will use to prove Lemma 3. Lastly, we provide a formal proof of the convergence time of log-linear learning in Theorem 1.
B.1 Proof of Theorem 1
We are now ready to proceed with the proof of Theorem 1.
Proof of Theorem 1.
By Lemma 3, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@ B.1 Proof of Theorem 1
Equation (2) in Lemma 1 provides an upper bound on the mixing time:
@@
where the third line follows from Lemma 2, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/2$, and the fact that $\Phi(\cdot)\in[0,1]$. This concludes the proof of Theorem 1. ∎
B.2 Proof of Lemma 2
Proof of Lemma 2.
Define the set $\mathcal{A}^{N}_{*}=\{a^{*}\in\mathcal{A}^{N}\>|\>a^{*}\in\operatorname*{arg\,max}_{a\in\mathcal{A}^{N}}\!\!\Phi(a)\}$ as the set of potential maximizers with cardinality $\!A^{N}_{*}\!=\!|\mathcal{A}^{N}_{*}|$. Then, the expected value of the potential function $\Phi(\cdot)$ over the stationary distribution $\mu$ of log-linear learning in (5) can be bounded as follows:
@@ B.2 Proof of Lemma 2
where we used that $\Phi(a)\leq 1$ for all $a\in\mathcal{A}^{N}$. ∎
B.3 Proof of Lemma 3
We first state the following lemma which will be used to prove Lemma 3.
Lemma 5.
@@ B.3 Proof of Lemma 3
First, note that the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ is aperiodic and irreducible with stationary distribution $\hat{\mu}(a)=1/A^{N}$. This can be verified by checking the detailed balance equations given by $\hat{\mu}(a)\hat{P}_{a,\tilde{a}}=\hat{\mu}(\tilde{a})\hat{P}_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. Next, we make use of Corollary $2.15$ in [26] to lower-bound $\rho(\hat{P})$ in terms of the log-Soblev constant of another Markov chain $\bar{X}_{t}$ with transition matrix $\bar{P}$ and stationary distribution $\bar{\mu}$. Namely,
@@
As the comparison Markov chain, we consider the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ with $\bar{X}_{t}=\prod_{i=1}^{N}\bar{X}_{i,t}$ on the state space $\mathbb{Z}_{K^{N}}=\prod_{i=1}^{N}\mathbb{Z}_{K}$. Here, each $\{\bar{X}_{i_{t}}\}_{t\in\mathbb{N}}$ is a simple random walk on the finite circle $\mathbb{Z}_{K}=\{1,\ldots,K\}$ with $K\geq 4$. The transition matrix and the stationary distribution of the simple random walk $\bar{X}_{i_{t}}$ are given by $\bar{P}_{i_{k,k\pm 1}}=\frac{1}{2}$ and $\bar{\mu}_{i}(k)=1/K$, respectively. Furthermore, the log-Soblev constant $\rho(\bar{P}_{i})$ is lower bounded by $\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25K^{2}}$ [10, Example 4.2]. Thus, by definition the product chain $\bar{X}_{t}$ [10, Sec. 2.5] has the following transition matrix:
@@
Furthermore, by Lemma $3.2$ in [10] the log-Soblev constant $\rho(\bar{P})$ of the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ is lower bounded by:
@@
which concludes the proof. ∎
Next, we proceed to prove Lemma 3.
Proof of Lemma 3.
The Markov chain $X_{t}$ with transition matrix $P$ defined in (7) is aperiodic and irreducible and thus a unique stationary distribution $\mu$ exists with $\mu^{t}=\mu^{0}P^{t}\rightarrow\mu$ for $t\rightarrow\infty$ from any initial distribution $\mu^{0}$. We define $\mu_{\max}$ and $\mu_{\min}$ as $\max_{a\in\mathcal{A}}\mu(a)\leq\mu_{\max}\leq 1$ and $\min_{a\in\mathcal{A}}\mu(a)\geq\mu_{\min}>0$, where $\mu_{\min}>0$ follows from the irreducibility of $X_{t}$.
@@
where we used that $P^{N}_{a,\tilde{a}}>0$ and $P_{a,a}>0$ for all $a,\tilde{a}\in\mathcal{A}^{N}$ as well as the identity $\mu(a)P^{*}_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$. It follows that $X^{*}_{t}$ is irreducible. Thus, for $X^{*}_{t}$ a unique stationary distribution $\mu^{*}$ exists. By Proposition $1.23$ in [15] the stationary distribution of $P^{*}$ is given by $\mu$ and thus the stationary distribution of $PP^{*}$ is given by $\mu$ since $\mu PP^{*}=\mu P^{*}=\mu$. Furthermore, the following holds for the transition matrix $PP^{*}$ :
@@
From Corollary $2.15$ in [10], it follows that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
@@
Next, we compare the Markov chain $X_{t}$ to the Markov chain $\hat{X}_{t}$ with transition matrix $\hat{P}$ specified in Equation (18) of Lemma 5. To this end, we compute $M$ and $C$ of the Markov chains $X_{t}$ and $\hat{X}_{t}$:
@@
Thus, by Corollary $2.15$ in [10] and by Lemma 5 the log-Soblev constant $\rho(P)$ can be lower-bounded by:

### v3

@@ 4.1 Reduced feedback
Log-linear learning requires players to access their utilities for all possible actions given the other players’ actions. Having such full-information feedback when action sets are large can be demanding. Binary log-linear learning (Arslan et al., 2007; Marden et al., 2007) alleviates this limitation by requiring two-point feedback, reducing the feedback needed by a factor $A$ per round. Now, we briefly review the binary-log-linear learning rule. Then, we derive the first finite-time convergence bound of binary log-linear learning to an $\epsilon$-efficient Nash equilibrium, showing that the deterioration in convergence time is a constant.
@@
$t\geq\frac{25N^{2}A^{5}}{2\pi^{2}}e^{4\beta}\left(\log\log A^{N}+\log\beta+2\log\frac{4}{\epsilon}\right)\approx\tilde{\mathcal{O}}\left(N^{2}A^{5}\left(\frac{A^{N}}{\epsilon}\right)^{\frac{N}{\max\{\epsilon,\Delta\}}}\right).$
Theorem 4.1 shows the first finite-time convergence guarantee for binary log-linear learning, we provide its detailed proof in Appendix C. It is remarkable that with significantly less feedback per round, binary log-linear achieves the same convergence speed as log-linear learning up to a factor of $8$ (Theorem 3.1). This raises the question of whether two-point feedback is sufficient for learning $\epsilon$-efficient Nash equilibria or whether the convergence time bounds we proved for log-linear learning are loose.
@@ 4.2 Perturbed log-linear learning
Classical log-linear learning relies on two limiting assumptions: 1) Players have access to their exact utilities. However, in real-world applications, the presence of noise is typical as uncertainties and hidden factors generate inexact measurements. 2) Players are rational. However, empirical evidence suggests that players have limited rationality and therefore may occasionally deviate from the log-linear learning rule in practical scenarios. Our next result generalizes Theorem 3.1 to the case where the log-linear learning rule is subject to small perturbations. As we will show, this generalization can address the two limitations above.
@@
We provide a proof in Appendix D.4. Compared to the result of (Zhang et al., 2023), we considerably improve the Lipschitz constant by using the mixing time bound based on log-Sobolev inequalities (Lemmas 2.2 and 3.4). In particular, (Zhang et al., 2023, Lemma 24) entails a Lipschitz constant $L=\tilde{\mathcal{O}}((e/p_{\min})^{N})$ while Lemma 4.2 implies that $L=\tilde{\mathcal{O}}(1/(\mu_{\min}p_{\min}^{3}))$. [^7] This can be seen by injecting the log-Sobolev bound of lemma 3.4 into our Lemma above. Now, we state the main result of this section.
Theorem 4.3.
@@
Theorem 4.3 shows that small perturbations of the log-linear learning rule do not compromise the convergence to an $\epsilon$-efficient NE. In particular, if the players follow a learning rule $P$ with $\|P-P_{\ell}\|_{2}=\mathcal{O}(\epsilon/(L\sqrt{A^{N}}))$ then they converge to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$. On the other hand, due to the unavailability of the stationary distribution of the perturbed learning rule, we suffer an extra factor of $(N/p_{\min})^{N}/N!$ in the convergence time guarantee compared to log-linear learning.
@@ 4.2.1 Corrupted utilities with additive noise
where $\xi_{i}(a_{i},a_{-i})\in[-\xi,\xi]$ is a bounded noise term. Alternatively, the noise could be assumed to be centered i.i.d. random variables with bounded variance (Leslie and Marden, 2011). Using Theorem 4.3, we hereafter prove that log-linear learning is robust to noisy feedback.
Corollary 4.4.
Consider the setting of Theorem 4.3 with noise-corrupted utilities as in Equation (9). If all players adhere to log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$ and $\xi\leq 1/(2\beta)$, then
@@
for $t=\mathcal{O}\left(N^{3/2}A^{3}e^{N+\beta(1+2\xi)(N+3)}\log\frac{1}{\epsilon^{2}}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
The proof is provided in Appendix D.2. Corollary 4.4 shows that log-linear learning with corrupted utilities converges to an $\epsilon$-efficient NE in time polynomial in $1/\epsilon$ if the corruption magnitude $\xi$ is sufficiently small. Our finite-time convergence result extends previous works on robust learning which provide asymptotic guarantees (Leslie and Marden, 2011; Lim and Shamma, 2013; Bravo and Mertikopoulos, 2017). The key to this result lies in showing that the transition matrix of the Markov chain induced by corrupted utilities is close to its corruption-free counterpart.
@@ 4.2.2 Log-linear learning mixed with uniform exploration
In the following, we assume players occasionally explore actions randomly. A modification of log-linear learning based on the fixed-share algorithm (Herbster and Warmuth, 1998) can reflect such a random behavior. In the so-called fixed-share log-linear learning, a player $i$ is randomly chosen and allowed to alter her action. Player $i$ samples her new action from the following distribution:
@@
The exploration parameter $\xi\in(0,1)$ determines how likely a player is to act randomly, where a value of $\xi=1$ corresponds to a uniform action sampling while $\xi=0$ corresponds to standard log-linear learning. For simplicity, we focus on the full-information case but fixed-share log-linear learning can easily be adapted to the binary setting. Note that this modification resembles the $\epsilon$-Hedge strategy (Heliou et al., 2017) in the expert advice literature, and under binary feedback, this modification resembles the Epx3.P strategy (Auer et al., 2002; Bubeck et al., 2012) in the multi-armed bandit literature. Here, the fixed share $\xi/A$ ensures a lower bound on the exploration.
Without explicit knowledge of the stationary distribution of this learning rule, we can apply Theorem 4.3 to deduce the following result.
Corollary 4.5.
Consider the setting of Theorem 4.3, where all players adhere to fixed-share log-linear learning with $\beta=\Omega\left(\frac{1}{\max\{\epsilon,\Delta\}}\log\frac{1}{\epsilon}\right)$. Then, for any $\epsilon\in(0,1)$ and initial distribution $\mu^{0}$ we have:
@@
for $t=\mathcal{O}\left(N^{3/2}A^{N+3}e^{\beta(N+3)}/(1-\xi)^{N+3}\right)$ with $L=\mathcal{\tilde{O}}\left(N^{2}A^{N+5}e^{\log(A^{N}/\epsilon)/\max\{\epsilon,\Delta\}}\right)$.
We provide a proof in Appendix D.3. Corollary 4.5 guarantees the convergence of fixed-share log-linear learning to an $\epsilon$-efficient Nash equilibrium in time polynomial in $1/\epsilon$ if the exploration parameter $\xi$ is sufficiently small. The key to this result is to show that the transition matrix of fixed-share log-linear learning is close to the transition matrix of the unperturbed learning rule in terms of the $\ell_{2}$ distance.
@@ Appendix A Markov chains and mixing times
In the following, we define the log-Sobolev constant and the spectral gap of a Markov chain. We then discuss the implications of Lemma 2.2 and provide a proof of the lemma.
Definition A.1 ((Diaconis and Saloff-Coste, 1996)).
@@
As mentioned in Section 2.2, using the log-Sobolev constant $\rho(PP^{*})$ can often significantly improve classical mixing time bounds based on the spectral gap $\lambda(PP^{*})$. Such classical bounds are of the form (Montenegro et al., 2006):
@@
where $C$ is some constant. Diaconis and Saloff-Coste (1996) in Lemma $3.1$ show that the log-Sobolev constant $\rho(PP^{*})$ is upper-bounded by the spectral gap $\lambda(PP^{*})$ as follows: $2\rho(PP^{*})\leq\lambda(PP^{*})$. Thus, roughly speaking if
@@
Proof A.2.
of Lemma 2.2 Let the relative entropy be defined as $D(\mu^{t}:\mu):=\sum_{a\in\mathcal{A}^{N}}\mu^{t}(a)\log\frac{\mu^{t}(a)}{\mu(a)}$. Then, for a Markov chain $\{X_{t}\}_{t\in\mathbb{N}}$ with irreducible transition matrix $P$ the relative entropy $D(\mu^{t}:\mu)$ decays at the following rate (Miclo, 1997):
@@
Note that $\rho(PP^{*})<1$ since $2\rho(PP^{*})\leq\lambda(PP^{*})$ by Lemma 3.1 in (Diaconis and Saloff-Coste, 1996) and for the spectral gap $\lambda(PP^{*})$ it is known that $\lambda(PP^{*})<1$ (Levin and Peres, 2017). To ensure that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon$, we derive the following lower bound on $t$:
@@ Appendix B Convergence of log-linear learning
We first prove Lemma 3.3 and then state a novel lemma which we will use to prove Lemma 3.4. Lastly, we provide a formal proof of the convergence time of log-linear learning in Theorem 3.1.
B.1 Proof of Theorem 3.1
We are now ready to proceed with the proof of Theorem 3.1.
Proof B.1.
of Theorem 3.1 By Lemma 3.4, the log-Sobolev constant $\rho(PP^{*})$ can be lower-bounded as:
@@
Equation (2) in Lemma 2.2 provides an upper bound on the mixing time:
@@
where the third line follows from Lemma 3.3, the fact that $\|\mu^{t}-\mu\|_{TV}\leq\epsilon/2$, and the fact that $\Phi(\cdot)\in[0,1]$. This concludes the proof of Theorem 3.1.
B.2 Proof of Lemma 3.3
Proof B.2.
[…]
@@
where we used that $\Phi(a)\leq 1$ for all $a\in\mathcal{A}^{N}$.
B.3 Proof of Lemma 3.4
We first state the following lemma which will be used to prove Lemma 3.4.
Lemma B.3.
@@
First, note that the Markov chain $\{\hat{X}_{t}\}_{t\in\mathbb{N}}$ is aperiodic and irreducible with stationary distribution $\hat{\mu}(a)=1/A^{N}$. This can be verified by checking the detailed balance equations given by $\hat{\mu}(a)\hat{P}_{a,\tilde{a}}=\hat{\mu}(\tilde{a})\hat{P}_{\tilde{a},a}$ for all $a,\tilde{a}\in A^{N}$. Next, we make use of Corollary $2.15$ in (Montenegro et al., 2006) to lower-bound $\rho(\hat{P})$ in terms of the log-Soblev constant of another Markov chain $\bar{X}_{t}$ with transition matrix $\bar{P}$ and stationary distribution $\bar{\mu}$. Namely,
@@
As the comparison Markov chain, we consider the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ with $\bar{X}_{t}=\prod_{i=1}^{N}\bar{X}_{i,t}$ on the state space $\mathbb{Z}_{K^{N}}=\prod_{i=1}^{N}\mathbb{Z}_{K}$. Here, each $\{\bar{X}_{i_{t}}\}_{t\in\mathbb{N}}$ is a simple random walk on the finite circle $\mathbb{Z}_{K}=\{1,\ldots,K\}$ with $K\geq 4$. The transition matrix and the stationary distribution of the simple random walk $\bar{X}_{i_{t}}$ are given by $\bar{P}_{i_{k,k\pm 1}}=\frac{1}{2}$ and $\bar{\mu}_{i}(k)=1/K$, respectively. Furthermore, the log-Soblev constant $\rho(\bar{P}_{i})$ is lower bounded by $\rho(\bar{P}_{i})\geq\frac{8\pi^{2}}{25K^{2}}$ (Diaconis and Saloff-Coste, 1996, Example 4.2). Thus, by definition the product chain $\bar{X}_{t}$ (Diaconis and Saloff-Coste, 1996, Sec. 2.5) has the following transition matrix:
@@
Furthermore, by Lemma $3.2$ in (Diaconis and Saloff-Coste, 1996) the log-Soblev constant $\rho(\bar{P})$ of the product chain $\{\bar{X}_{t}\}_{t\in\mathbb{N}}$ is lower bounded by:
@@
which concludes the proof.
Next, we proceed to prove Lemma 3.4.
Proof B.5.
of Lemma 3.4 The Markov chain $X_{t}$ with transition matrix $P$ defined in (7) is aperiodic and irreducible and thus a unique stationary distribution $\mu$ exists with $\mu^{t}=\mu^{0}P^{t}\rightarrow\mu$ for $t\rightarrow\infty$ from any initial distribution $\mu^{0}$. We define $\mu_{\max}$ and $\mu_{\min}$ as $\max_{a\in\mathcal{A}}\mu(a)\leq\mu_{\max}\leq 1$ and $\min_{a\in\mathcal{A}}\mu(a)\geq\mu_{\min}>0$, where $\mu_{\min}>0$ follows from the irreducibility of $X_{t}$.
@@
where we used that $P^{N}_{a,\tilde{a}}>0$ and $P_{a,a}>0$ for all $a,\tilde{a}\in\mathcal{A}^{N}$ as well as the identity $\mu(a)P^{*}_{a,\tilde{a}}=\mu(\tilde{a})P_{\tilde{a},a}$. It follows that $X^{*}_{t}$ is irreducible. Thus, for $X^{*}_{t}$ a unique stationary distribution $\mu^{*}$ exists. By Proposition $1.23$ in (Levin and Peres, 2017) the stationary distribution of $P^{*}$ is given by $\mu$ and thus the stationary distribution of $PP^{*}$ is given by $\mu$ since $\mu PP^{*}=\mu P^{*}=\mu$. Furthermore, the following holds for the transition matrix $PP^{*}$ :
@@
From Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996), it follows that the log-Soblev constant $\rho(PP^{*})$ is lower-bounded by:
@@
Next, we compare the Markov chain $X_{t}$ to the Markov chain $\hat{X}_{t}$ with transition matrix $\hat{P}$ specified in Equation (18) of Lemma B.3. To this end, we compute $M$ and $C$ of the Markov chains $X_{t}$ and $\hat{X}_{t}$:
@@
Thus, by Corollary $2.15$ in (Diaconis and Saloff-Coste, 1996) and by Lemma B.3 the log-Soblev constant $\rho(P)$ can be lower-bounded by: